OpenStudy (anonymous):

DEFG is a rectangle DF=5x-5, and EG=x+1. Find the values of x and the length of each diagonal a.)x=4;DF=13;EG=13 b.)x=4;DF=15;EG=18 c.)x=4;DF=15;EG=15 d.)x=2;DF=13;EG=13

OpenStudy (anonymous):

@BunnyBree are you available at all ??

OpenStudy (anonymous):

@Timmyk14 are you available??

OpenStudy (anonymous):

I need help with this :)

OpenStudy (timmyk14):

Okay, give me a moment to look over the question.

OpenStudy (anonymous):

alrighty :)

OpenStudy (anonymous):

Are you sure it's EG = x + 1?

OpenStudy (anonymous):

its EG=x+11 sorry

OpenStudy (anonymous):

Okay. So the answer would be C.

OpenStudy (timmyk14):

This one, like the other, is basically asking you to plug in the value of x given and solve. So we start with answer choice A and work our way to the correct answer. Answer choice A says that the value of x is equal to 4. So we use the equation they gave us (DF = 5x -5 and EG = x + 1) to solve. So... DF - 5(4) - 5 DF = 20 - 5 DF = 15 EG = 4 + 1 EG = 5 Instantly we notice that this answer choice is wrong. This is because it says that DF is equal to 13, where we just proved that it's not. So onto the next answer choice..B! Okay, @BunnyBree is right. None of the answer choices could be correct if EG = x+ 1. Double check to see if there isn't supposed to a coefficient to x.

OpenStudy (timmyk14):

Well, we are supposed to explain to you why it is that answer choice, not just give you the answer. But do you understand how she go that answer? It's like we did in the last question. And exactly the same to what I did above.

OpenStudy (anonymous):

DF = 5(4) - 5 DF = 20 - 5 DF = 15 EG = 4 + 11 EG = 15 X = 4, DF = 15, EG = 15. So, C. Do you understand now? (: